OpenStudy (anonymous):

simplify mn^-4 ------- p^0q^-2

zepdrix (zepdrix):

So by simplify I assume they want `no negative exponents in your result, yes?

zepdrix (zepdrix):

\(\Large\rm \dfrac{m^{-4}}{p^0q^{-2}}\) Start with the p first. Do you emember what happens when you raise something to the 0 power?

OpenStudy (jdoe0001):

\(\bf \cfrac{mn^{-4}}{p^0q^{-2}}\implies \cfrac{m}{1}\cdot \cfrac{1}{p^0}\cdot \cfrac{n^{-4}}{1}\cdot \cfrac{1}{q^{-2}}\implies ? \\ \quad \\ \textit{keep in mind that }a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies \large a^{{\color{red} n}} \)